OpenStudy (anonymous):

A boy walked 10 kilometers. He returned in a car at the rate of 30 kilometers an hour. He was gone for 4 hours. What was his rate while walking?

OpenStudy (anonymous):

It took him 20 minutes to drive back. 10 km / 30 km/hr = 1/3 hour = 20 min That means it took him 3 2/3 hours to walk 10 km 3 2/3 = 11/3 hours 10/(11/3) = 30/11 = 2.73 km/hr